1. Calculate the percent composition by mass of ​water in the compound CaSO​4​ • 6 H​2O​ ​ [​GFM = 244].


2​.​ Calculate the percent composition by mass of ​water in the hydrate CaCl​2 ​ • 2H​2O​ .


3. ​Calculate the percent composition by mass of water​ in the hydrate ​CuSO​4​ • 5 H​2O​ ?


4. Calculate the percent composition by mass of water​ in the hydrate CaO​ ​ • 4H​2O​ [GFM = 128]


5. ​ ​What is the percent mass concentration of 45 g of a solution containing 9.0 g of KClO​3


6 ​ A single 500.0 gram serving of one sports drink contains 0.055 grams of sodium ions. Calculate the concentration of sodium ions in this sports drink, expressed as percent by mass.


1
Expert's answer
2021-02-01T03:43:14-0500
"\omega H2O=M(H2O)/M (CaSO4*6H2O)*100 %"
"\omega H2O=6*18/244*100=44.26%"

2.

"\omega H2O=2*18/147*100=24.49 %"

3.

"\omega H2O=5*18/250*100=36 %"

4.

"\omega H2O=4*18/128*100=56.25 %7"

5.

"\omega KClO3=m KClO3/m*100 = 9/45 * 100=20 %"

6.

"\omega K+=m(K+)/m*100=0.055/500*100=0.011 %"